A square plot of land is covered with cement costing $8/m^2 and a fence surrounding it costs $10/m. If the total cost is $8400, find the area of the plot
Answer by Alan3354(69443)

A square plot of land is covered with cement costing $8/m^2 and a fence surrounding it costs $10/m. If the total cost is $8400, find the area of the plot
-----------------
side length = s
----
Cost = 8s^2 + 4*10s = 8400
8s^2 + 40s - 8400 = 0
s^2 + 5s - 1050 = 0
(s-30)*(s+35) = 0
s = 30 meters
Area = 900 sq meters